OpenStudy (anonymous):

a= (2011^2 + 2012)^(1/2) if f(x)=[x],[x] denotes greatest integer fn. then value of f(a) is?

OpenStudy (anonymous):

a= \[\sqrt{2011^{2}+2012}\]

ganeshie8 (ganeshie8):

\(\large \sqrt{2011^2 + 2011 + 1}\)

ganeshie8 (ganeshie8):

complete the square, and mess wid the property : \((a+1)^2 - a^2 = 2a + 1\)

ganeshie8 (ganeshie8):

\(\large \sqrt{(2011 + \frac{1}{2})^2 + \frac{3}{4}}\)

ganeshie8 (ganeshie8):

\(\large 2011 \lt \sqrt{(2011 + \frac{1}{2})^2 + \frac{3}{4}} \lt 2012\)

OpenStudy (anonymous):

what is the value of a @ganeshie8

ganeshie8 (ganeshie8):

^^look at above it lies between 2011 and 2012

ganeshie8 (ganeshie8):

you dont need the exact value to evaluate greatest integer function

OpenStudy (anonymous):

not that.. the a in ur eq @ganeshie8

ganeshie8 (ganeshie8):

\( a = \sqrt{2011^{2}+2012} = \sqrt{2011^2 + 2011 + 1} \\ = \sqrt{2011^2 + 2\times \frac{1}{2} \times 2011 + \frac{1}{4} + \frac{3}{4}} \\ = \sqrt{(2011 + \frac{1}{2})^2 + \frac{3}{4}}\)

OpenStudy (anonymous):

@ganeshie8 how can you assure it is leass than 2012 ?

ganeshie8 (ganeshie8):

use below : \((a+1)^2 - a^2 = 2a + 1 \)

ganeshie8 (ganeshie8):

the difference between any two consecutive squares is \(2a + 1\)

ganeshie8 (ganeshie8):

since 3/4 is no way near to 2011, you can be confident that it will not go near 2012

ganeshie8 (ganeshie8):

\((2011+1)^2 - 2011^2 = 2*2011 + 1 \)

ganeshie8 (ganeshie8):

to jump from *square* of 2011 to 2012, u need an extra of 2*2011 + 1 but ur 3/4 is very less - so it will not result in approaching 2012

OpenStudy (anonymous):

 @ganeshie8 it is plus not minus